A mixture of methane and carbon dioxide gases, at a total pressure of 648 mm Hg, contains 2.03 grams of methane and 6.40 grams of carbon dioxide. What is the partial pressure of each gas in the mixture? PCH4 =

PCO2 =


1
Expert's answer
2022-03-28T21:29:05-0400

Ratio 2.03:6.40

PCH4= (2.03/8.43) × 648= 156mmHg

PCO2= 648 - 156= 492mmHg